OpenStudy (anonymous):

You invest $1,600 in an account that pays an interest 6.5%, compound continuously. Calculate your balance after 6 years. Round to the nearest hundredth.

OpenStudy (anonymous):

Can anyone show me step by step on how to do this, like if you were explaining this to a 5 year old because I am clueless to Exponential and Logarithmics!

OpenStudy (wolf1728):

6.5 % interest compounded continuously = (e^r) -1 = (2.718181828^ .065) -1 =1.0671590244 -1 =.0671590244 =6.71590244 % compounded continuously

OpenStudy (anonymous):

But how did you get (2.718181828^ .065) What did you do to get those numbers?

OpenStudy (wolf1728):

Okay I went to this web page http://www.1728.org/rate.htm and used the 3rd formula on that page 'e' is the base of natural logarithms 2.718281828 Yes, I know that's a lot for a 5 year old but I'll try to explain what you need to know.

OpenStudy (anonymous):

What do you mean by base of natural logarithms? Sort of like how Pi = 3.14?

OpenStudy (wolf1728):

Compounded interest is tough enough to explain but what intructor gives you a CONTINOUSLY compounded interest problem?

OpenStudy (wolf1728):

Let's just act if this were a compounded interest problem

OpenStudy (wolf1728):

The formula for the total of a compounded interest account is TOTAL = Principal * (1 + rate)^years (Page explaining compound interest is here: http://www.1728.org/compint2.htm)

OpenStudy (anonymous):

The question does show this formula but, idk how to apply it and then solve.

OpenStudy (wolf1728):

1,600.00 invested 6 years at 6.5% TOTAL = principal * (1+rate)^years TOTAL = 1,600.00 * (1.065)^6 TOTAL = 1,600.00 * (1.4591422965) TOTAL = 2,334.63 What formula did they tell you?

OpenStudy (anonymous):

A = Pe^rt ??

OpenStudy (wolf1728):

Okay TOTAL = Principal * e^(rate * time) TOTAL = 1,600 * 2.718281828 ^ (.065 * 6) TOTAL = 1,600 * 1.4769807938 TOTAL = 2,363.17

OpenStudy (anonymous):

Wait, how do I calculate 2.718281828 ^ (.065 * 6) to get 1.4769807938? Like 2.7182 0.65, I'm still doing it wrong. :(

OpenStudy (wolf1728):

I guess I first should have shown you this calculation .065 * 6 = 0.39 Then next we calculate e^.39 or 2.718281828 ^ .39

OpenStudy (anonymous):

I got 1.06012991292 ?? instead of 1.4769807938

OpenStudy (wolf1728):

You are raising a number to a power on your calculator, correct?

OpenStudy (anonymous):

I'm using the standard calculator in my laptop, Sorry I'm so dumb, online math class is so difficult for me and I miss having a teacher face to face.

OpenStudy (wolf1728):

Yeah, even that formula you posted A = Pe^rt doesn't really explain anything does it?

OpenStudy (anonymous):

This one is different concerning this formula A = p(1 + r)^t Tina Invests $1,000 with interest rate 6.75% How many years will it take for the account to reach $18,600 ?

OpenStudy (anonymous):

Hopefully I can understand these ones better than A = Pe^rt

OpenStudy (wolf1728):

Hi LadyLillyz I'm back

OpenStudy (anonymous):

I'm still here!

OpenStudy (wolf1728):

First, let's see if you can raise a number to a power. Do you know how to calculate this? 2^3

OpenStudy (wolf1728):

Yes, you could just calculate 2 *2 * 2 but can you do it with your calculator?

OpenStudy (anonymous):

Negative.

OpenStudy (anonymous):

Any other possible sources via online calculators?

OpenStudy (wolf1728):

You don't know how to raise a number to a power? You'll DEFINITELY need to know that in order to do these problems. Let me see about online calculators.

OpenStudy (wolf1728):

Here's one right here: http://www.1728.org/calc.htm

OpenStudy (anonymous):

I do know how to raise a number to a power, just not the complicated one like we did previously "2.718281828 ^ .39"

OpenStudy (wolf1728):

Okay go here but fore you do http://www.1728.org/calc.htm try raising 2 to the third power 2^3 = ?

OpenStudy (anonymous):

It is 8

OpenStudy (wolf1728):

You used the calculator right?

OpenStudy (anonymous):

I used the calculator after raising it like you asked.

OpenStudy (wolf1728):

Okay - now to get more accuracy out of that calculator, go to the bottom of that page and change the 5 to a 20

OpenStudy (anonymous):

Objective done.

OpenStudy (wolf1728):

Now just to see if it changed try 2^.5

OpenStudy (wolf1728):

You should get 1.4142135623730951

OpenStudy (anonymous):

1.4142135623730951, 100% match

OpenStudy (wolf1728):

All right we are set to do some continuously compounded interest problems!!!

OpenStudy (wolf1728):

Okay why not try 2.718281828 ^ .39

OpenStudy (anonymous):

1.4769807937853678,

OpenStudy (wolf1728):

Well, now you know how to raise a number to a power. Now if you look way back to where that had to be calculated, now you know where I got that number. If not I'll just type it all again.

OpenStudy (wolf1728):

Here's the original problem: TOTAL = Principal * e^(rate * time) TOTAL = 1,600 * 2.718281828 ^ (.065 * 6) TOTAL = 1,600 * 2.718281828 ^ (.039) TOTAL = 1,600 * 1.4769807938 TOTAL = 2,363.17

OpenStudy (anonymous):

Okay, it's getting easier now, but the other question earlier, with a different formula and a different question..Tina Invests $1,000 with interest rate 6.75% How many years will it take for the account to reach $18,600 ? It doesn't state x amount of money after years but instead, the opposite?

OpenStudy (wolf1728):

1,000 rate =6.75 when does it reach $18,600? Is this continuously compounded?

OpenStudy (anonymous):

It gives me a formula A = P(1 + r)^t

OpenStudy (wolf1728):

That's what your book says? Okay well that is regular compound interest. But that has to be solved for 't'. If you can't find that formula I'll show you how to do it.

OpenStudy (anonymous):

I don't want to force you to give me a full lecture and waste your time, I think 1 more solid example and good explaination I can do other problems independently, but I do come across another speed bump, I will notify you.

OpenStudy (wolf1728):

TOTAL = principal * (1 + rate)^t TOTAL / principal = (1 + rate)^t take logarithms of both sides log (TOTAL / Principal) = time * log (1 + rate) time = log (TOTAL / Principal) / log (1 + rate) time = log (18,600 / 1,000) / log (1.065) time = 44.751767801 (Okay I left out all the log calculations)

OpenStudy (wolf1728):

I'd say THIS problem was a speed bump unless you think you could have solved the formula for time. LOL